OpenStudy (2yoututu11):

What is the volume of the cylinder? A. 126 ft3 B. 252 ft3 C. 882 ft3 D. 2268 ft3

OpenStudy (anonymous):

V = pi*r^2h

OpenStudy (atlas):

Volume of a cylinder is given by pi*r*r*h; Here, r is the radius of the base of the cylinder and h is the height

OpenStudy (anonymous):

what is hieght i can't see what number it is?

OpenStudy (2yoututu11):

there is no height

OpenStudy (anonymous):

@ZhangYan , try control + to make your screen bigger.

OpenStudy (anonymous):

height is 18 ft

OpenStudy (anonymous):

i can't dpalnc not on my computer sorry

OpenStudy (2yoututu11):

what is the browser are u on

OpenStudy (anonymous):

A = 3.14 * 7^2* 18

OpenStudy (anonymous):

can you solve it

OpenStudy (anonymous):

oh and the browser is firefox

OpenStudy (2yoututu11):

ohhh I use chrome have u ever used it

OpenStudy (anonymous):

no just firefox...

OpenStudy (anonymous):

can you solve this? A = 3.14 * 7^2* 18

OpenStudy (2yoututu11):

u should use chome it is way better u can choose your backround and it saves your websites so u won't have to remember all the websites u want to go to ..

OpenStudy (2yoututu11):

ohh and yes I can solve it

OpenStudy (2yoututu11):

thank you

OpenStudy (anonymous):

nah i am fine i have INternet Exporer and firefox so i am fine :) but thanks

OpenStudy (anonymous):

np :D